Solve the quadratic equation by completing the square:
4(x)^2 – 12(x) – 55 = 0
4(x)^2 – 12(x) = 55
4(x)^2 – 3(x) = 55
(x)^2 – 3(x) = 55/4

I know I have to add (b/2)^2 to both sides of the equation, which I calculated as 9/4...or do I keep this as (3/2)^2?
(x^2 – 3x + 9/4) = (55/4 + 9/4)
(x^2 – 3x + 9/4) = 16
And this is where I get lost. There seem to be so many different ways this is done and I can't figure out what my goal is. I've been trying to figure this out for a while and could use a little guidance. Thanks for your time :-)

Answer by edjones(8007)

Solve the quadratic equation by completing the square:
4x^2–12x–55=0
4x^2–12x=55
x^2–3x= 55/4 The line above this was incorrect and I removed it. This is correct.
x^2-3x+9/4=9/4 + 55/4
(x - 3/2)^2 =16 Factor left side
x - 3/2 = +-4 sqrt of each side
x=1.5 +-4 [3/2=1.5]
x=5.5, x=-2.5
.
You did the hard part.
.
Ed